A football is kicked at 45 degrees and travels 82 meters before hitting the ground.

A) What was its initial velocity?
B) How long was it in the air?
C) How high did it go?
Answer #1
In the horizontal...
vh= vi*cos45
82=vi*cos45*timeinair.
solve for time in air.
In the vertical:
hf=hi+vi*sin45*t - 4.9t^2
0=0+vi*sin45t-4.9t^2
solve for t, the time in air.
set that expression equal to the first timeinair expression, then solve for vi
then, get a numberfor time in air from vi.

Then finally, at the top, vy=0
vy=0=vi*sin45-gt
solve for t, the time to top.
htop= vi*sin45*t-1/2 g t^2
answered by: Franny